### v1

Abstract
This work presents an approach to improve the precision of polarized photon-photon collisions simulation implemented in the SANCphot package. The basic linear Compton approximation of the incoming photon spectrum is extended to a general energy distribution and a realistic description of circular or linear polarizations as expected to be seen at photon-photon colliders.
keywords:
Photon collider, Photon-photon collisions, Polarized photons, Perturbation theory, NLO calculations, Standard Model, Electroweak interaction, QED, Monte Carlo integration
1 Introduction
Photon colliders, envisioned as potential future high-energy physics research instruments, offer a unique opportunity to explore electroweak and beyond-the-Standard-Model (BSM) physics in unprecedented precision. While the phenomena of photon-photon collisions were previously observed as a side effect in electron-positron and heavy ion collisions, a dedicated machine would open up new avenues for precision electroweak physics. An overview presented in [1] covers a wide range of topics dedicated to the development of a photon-photon collider, including physics program, photon beam generation, and effects in the interaction region. Theoretical calculations and simulation codes accompany these developments, gradually improving our understanding and correcting expectations from these intriguing projects.
In this paper, we investigate the realistic simulation of polarized photon-photon collisions at photon colliders by coupling the detailed photon beam simulation and the theoretical calculations for the $\gamma\gamma$ processes. A precise description of the $\gamma\gamma$ processes that considers initial photon polarization is useful in a couple of aspects. When studied in angular distributions, polarization effects of the Z-boson in processes $\gamma\gamma\rightarrow\gamma Z$ and $\gamma\gamma\rightarrow ZZ$ provide sensitivity to potential new physics beyond the Standard Model. Anomalous quartic gauge couplings and extra dimensions are another kind of phenomena, where accurate simulation of polarized photon-photon collisions can improve background estimation and lead to better constraints [2] in these domains.
Previously, several remarkable works have been devoted to the study of $\gamma\gamma$ processes, in particular, light-by-light scattering (LbL), taking into account radiative corrections [3, 4, 5, 6, 7]. The process of $\gamma\gamma\to\gamma Z$ was revized in [8]. A calculation of ZZ production cross-section components ($\sigma_{0},\sigma_{22},$ etc.) was presented and available as a public code [9]. Processes $\gamma\gamma\rightarrow\gamma\gamma$ and $\gamma\gamma\rightarrow\gamma Z$ with polarized incoming beams at CLIC are investigated in terms of anomalous couplings [10, 11].
The paper is composed as follows. A short preliminary word on the current status of the photon-photon collider is given in the Introduction 1. We revise how photon-photon collisions are treated in the SANCphot code in Section 2. Then, in Section 3, we overview challenges occurring when trying to give a precise description of photon spectra generated at the photon-photon colliders by the Compton scattering. Our method of supplying a realistic photon spectrum to the SANCphot calculation is described in Section 4, and numeric results of this work are presented in Section 5. The summary in Section 6 gives a brief overview of our studies.
2 Photon-photon collisions in SANC
The cross section of photon-photon collisions implemented in the SANCphot tool [12] follows a recipe described in [13] using FORTRAN modules introduced in [14, 15, 16]. The total $\gamma\gamma$ process cross section is represented as a convolution of its components in different helicity combinations $d\bar{\sigma}_{ij}/d\cos\vartheta^{*}$ and averaged Stokes parameters defining the density matrix of the back-scattered photon $\langle\xi_{i}\xi_{j}^{\prime}\rangle$:
$\begin{split}{\frac{d\sigma}{d\tau d\cos\vartheta^{*}}}=&{\frac{d\bar{L}_{\gamma\gamma}}{d\tau}}\Bigg{\{}{\frac{d\bar{\sigma}_{0}}{d\cos\vartheta^{*}}}+\langle\xi_{2}\xi_{2}^{\prime}\rangle{\frac{d\bar{\sigma}_{22}}{d\cos\vartheta^{*}}}\\ &+[\langle\xi_{3}\rangle\cos 2\phi+\langle\xi_{3}^{\prime}\rangle\cos 2\phi^{\prime}]{\frac{d\bar{\sigma}_{3}}{d\cos\vartheta^{*}}}\\ &+\langle\xi_{3}\xi_{3}^{\prime}\rangle[{\frac{d\bar{\sigma}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi+\phi^{\prime})\\ &+{\frac{d\bar{\sigma}^{\prime}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi-\phi^{\prime})]\\ &+[\langle\xi_{2}\xi_{3}^{\prime}\rangle\sin 2\phi^{\prime}\\ &-\langle\xi_{3}\xi_{2}^{\prime}\rangle\sin 2\phi]{\frac{d\bar{\sigma}_{23}}{d\cos\vartheta^{*}}}\Bigg{\}}\ ,\end{split}$ (1)
In this definition, the $\vartheta^{*}$ is the scattering angle of $\gamma\gamma$ collision products measured in the rest frame of the incoming particles, and the $\tau\equiv s_{\gamma\gamma}/s_{ee}$ — a fraction of the electron beam’s energy transferred to the back-scattered photons. Further, $d\bar{L}_{\gamma\gamma}/d\tau$ denotes the photon-photon luminosity per unit $e^{+}e^{-}$ flux, considering the linear collider operating in $\gamma\gamma$ mode, and the Stocks parameters $\xi_{2},\xi_{3}$ are defined via the photon and electron polarization functions as described in [17][18]. The $d\bar{L}_{\gamma\gamma}/d\tau$ parameter, consequently, depends on the energy distribution of the back-scattered photons and can be evaluated analytically using the linear Compton approximation (LCA).
3 Photon spectrum in photon colliders
The linear Compton approximation has several drawbacks when utilized in the simulation of the $\gamma\gamma$ collisions in the photon-photon collider [19]. Most of the inaccuracy is attributed to the process of electron beam conversion to a high-energy photonic beam. Here, the oncoming laser beam has high but nonuniform intensity, and the electron bunches have distorted geometry, nonuniform electron density, and a strong electromagnetic field. More difficulties are added by the fact that the particle momenta are not perfectly aligned with the beam axis and the presence of an external magnetic field at the interaction point.
Several works have been attributed to improving the spectrum of the photons generated via electron energy transfer. To our knowledge, the said issues are best addressed in the photon collider simulation code CAIN [20]. The substantial improvements are achieved by introducing full treatment of nonlinear and polarized Compton scattering, a detailed description of the space-time structure of the electron and laser beams, complete kinematics, including angular distributions, and interplay with beamstrahlung and other background processes. Other efforts to improve photon spectrum approximation are presented in works [21, 22] which contribute to a detailed description of nonlinear effects and a complete description of the photon polarization in Compton back-scattering.
Public codes suitable for photon spectrum simulation include CompAZ [23] and CIRCE1 [24], both based on the creation of explicitly parameterized distributions. These tools are focused on approximating the high-energy regions of the photon spectra. CIRCE2 [25] features correlated double-beam spectra based on tabulated energy distributions built from the event records generated in advance.
4 Coupling CAIN and SANCphot
In order to supply the realistic photon spectra obtained with the CAIN simulation, including the polarization information, to the Monte Carlo integration routines implemented in SANCphot, we use an approach similar to CIRCE2 [25]. The event records, generated with a realistic laser and oncoming electron beam parameters, along with a non-linear Compton process expressions, are aggregated into piecewise continuous functions, representing approximate distributions of spectra and linear and circular polarizations.
$f_{L}(x)=\sum_{i=0}^{n}\bar{L}_{i}\chi_{i}(x),\;\chi_{i}(x)=\begin{cases}1\;\mathrm{if}\;\left\llbracket x/n\right\rrbracket=1\\ 0\;\mathrm{if}\;\left\llbracket x/n\right\rrbracket\neq 1\\ \end{cases}$ (2)
where $f_{L}$ is a piecewise approximation of the luminosity, $\chi_{i}(x)$ is an indicator function, $L_{i}$ is the aggregated $L_{\gamma\gamma}$ value in bin $i$. Similar definitions are used for approximations of the Stokes parameters: $f_{\xi_{2}}(x)$ and $f_{\xi_{3}}(x)$. The cubic interpolation of these functions serves as an input spectrum to SANCphot’s integration procedures in the same manner as analytic expressions of linear Compton.
As an example of a realistic photon spectrum simulation, the parameters of the CLIC accelerator was chosen to configure the CAIN code: number of electrons per bunch $N=4\times 10^{9}$, the RMS (root-mean-square) bunch length in the longitudinal direction $\sigma_{z}=0.05$ mm, normalized emittance in the y direction $\gamma\varepsilon_{y}=1\times 10^{-5}\mathrm{m}\cdot\mathrm{rad}$ and in x direction $1.9\times 10^{-6}$ and $1.5\times 10^{-6}\mathrm{m}\cdot\mathrm{rad}$ for 500 and 1000 GeV respectively, beta functions $\beta_{x,y}=2/0.1$ mm and the transverse beam sizes $\sigma_{x,y}=88/4.5$ nm for 500 GeV and $55/3.2$ nm for 1000 GeV.
The polarizations of the back-scattered photons in CAIN were controlled via Stokes parameters $(\xi_{1},\xi_{2},\xi_{3})$. Simulation of non-linear Compton scattering is controlled via the number of absorbed laser photons $n_{ph}$ parameter in CAIN. However, the support for electron transverse polarization is limited only to linear Compton approximation, and for values $n_{ph}>0$ the laser photon polarization is narrowed to the circular one (i.e. $\xi_{1}=\xi_{3}=0,\,\xi_{2}=\pm 1$). The difference between $n_{ph}=0$ and $n_{ph}=1$ setting is that the former uses linear Compton formulas, which are inapplicable in strong laser fields. In the CAIN simulation, the strength of a laser electromagnetic field is characterized by a parameter $\xi=\frac{\lambda_{L}}{m}\sqrt{\mu_{0}cP}$, where $\lambda_{L}$ is the laser wavelength, m is the electron mass, $c$ is the speed of light, $\mu_{0}=4\pi\times 10^{-7}$, and P is the laser power density. At $\xi<O(1)$, the non-linear Compton expression requires expansion in terms of $n_{ph}$ to provide a better description.
Figure 1: A comparison of photon energy spectra $dL/dx$ and cross section polarizing coefficients $d\langle\xi 2\rangle/dx$ and $d\langle\xi 3\rangle/dx$ depending on the fraction of transferred energy $x$. Solid lines and dashed lines correspond to linear and circular laser polarizations.
In the present study, we use two distinct polarization configurations:
  • linear : $P_{e}=P_{e}^{\prime}=0,P_{\gamma}=P_{\gamma}^{\prime}=0,P_{t}=P_{t}^{\prime}=1,\phi=\pi/2$
  • circular : $P_{e}=P_{e}^{\prime}=0.8,P_{\gamma}=P_{\gamma}^{\prime}=-1,P_{t}=P_{t}^{\prime}=0$
Here $P_{e}$ is an electron longitudinal polarization, $P_{\gamma}$ describes the average helicity of the laser photon, $P_{t}$ denotes maximum average transverse polarization along a direction determined by the azimuthal angle $\phi$. This angle is determined with respect to the back-scattered photon momentum.
Figure 1 shows photon energy spectra and polarization at different levels of approximation: LCA calculated with SANCphot (black lines), linear and non-linear Compton approximations calculated by CAIN with $n_{ph}=$ 0 (red), 1 (blue), and 2 (green) with realistic beam geometry. The CAIN spectra show a significant shift towards lower energy regions due to multiple electron-photon interactions and spatial beam effects.
5 Numerical results
The numerical validation was performed in the $\alpha(0)$ EW scheme using the setup given below. For ease of validation, we keep the electroweak parameter values the same as in [12]:
$\alpha=1/137.035990996$, $G_{F}=1.13024\times 10^{-5}\,\mathrm{GeV}^{-2}$, $M_{W}=80.45149$ GeV, $M_{Z}=91.18670$ GeV, $M_{H}=125$ GeV, $m_{e}=0.51099907$ MeV, $m_{\mu}=0.105658389$ GeV, $m_{\tau}=1.77705$ GeV, $m_{u}=0.062$ GeV, $m_{d}=0.083$ GeV, $m_{c}=1.5$ GeV, $m_{s}=0.215$ GeV, $m_{t}=173.8$ GeV, $m_{b}=4.7$ GeV.
The simulation of gamma-gamma collisions was performed as if the photons were produced in the $ee$ collider framework. Initial photon spectra were approximated using several methods: (i) analytic linear Compton approximation implemented in SANCphot, (ii) photon energy distribution representing linear Compton spectrum, (iii) realistic spectra obtained from CAIN photon collider simulation package. The initial electron energy was picked from 500 GeV and 1 TeV, and the laser beam energy irradiating the electrons was set to 1.26120984 eV, assuming that it provides an optimal conversion ratio for the back-scattered photons with parameter $x_{0}=4.83$ (at electron beam energy 500 GeV). The analytic expressions for the 4-boson interaction diverge at large scattering angles, making the numeric calculation unstable. Therefore, the phase space is constrained by $30^{\circ}<\theta<150^{\circ}$. Also the final state invariant mass for the processes $\gamma\gamma\rightarrow\gamma\gamma,\,\gamma\gamma\rightarrow\gamma Z,\,\gamma\gamma\rightarrow ZZ$ was constrained to be above 20, 100 and 200 GeV correspondingly to ignore kinematically forbidden regions.
LCA SANCphot SANCphot CAIN
spectrum analytic piecewise piecewise
linear 8.2019(4) 8.2023(4) 8.4657(4)
circular 8.8809(4) 8.8811(4) 9.3032(5)
Table 1: Integrated cross sections $\sigma(\gamma\gamma\rightarrow\gamma\gamma)$ [fb] for linear Compton approximation spectra.
We perform several checks to validate the simulation of photon-photon collisions with arbitrary initial photon spectra and polarization distributions. Firstly, we compare the total cross sections and kinematic distributions in azimuthal angle, invariant mass, and transverse momentum computed using analytic expressions for linear Compton approximation and those calculated via pre-generated LCA photon spectra for transverse and circular photon polarizations shown in Figure 2. In these cases, the photon conversion point coincides with the central interaction point. Second, we compare with the cross sections based on the spectra obtained with the photon-photon collision simulation by CAIN to validate against established simulation package and justify its further usage for realistic spectra simulation. CAIN simulates the realistic beam shapes and angular distributions, therefore, despite its settings being set to linear Compton $(n_{ph}=0)$, the spatial effects produce visible differences from the pure Compton scattering.
Figure 2: The kinematic distributions for $\gamma\gamma\rightarrow\gamma\gamma$ process at 500GeV with linear Compton approximation. Upper plots are based on the analytic expressions from SANCphot. Lower plots give the ratios of distributions built using piecewise approximation of SANCphot (red) and CAIN (blue) spectra to the analytic values.
In addition, Table 1 shows numeric comparison of three different sources of photon spectra, when using the analytic version from SANCphot, piecewise spectrum description of the analytic expression, and LCA obtained with CAIN minimizing beam geometry effects (i.e., using a very large beam, very low laser intensity, and very low number of particles per bunch). We see that the piecewise description of pure LCA spectrum is very close to the analytic one.
For a more precise photon beam simulation in the beam parameters similar to the CLIC accelerator, we consider the invariant mass and transverse momentum distributions plotted on the Figures 4 and 5. The plots show differential cross sections calculated using analytic expressions for linear Compton approximation (black lines), linear Compton with realistic beam geometry simulated in CAIN (red lines), and taking into account non-linear Compton effects together with a realistic beam geometry simulated in CAIN for $n_{ph}=1$ (blue) and $n_{ph}=2$ (green). The solid and the dashed lines correspond to circular and linear photon polarization, respectively.
Figure 3: Final state $\cos\theta$ kinematic distributions for electron beam energy $\sqrt{s_{ee}}=500$ GeV and various polarizations (see text).
Figure 4: Final state invariant mass kinematic distributions for different electron beam energies $\sqrt{s_{ee}}=500$ GeV (upper row) and 1000 GeV (lower row) and various polarizations (see text).
In the invariant mass distributions, we observe a well-known energy-dampening effect of taking into account realistic beam geometry. The shift to the less energetic area of spectra is mostly visible for the $\gamma\gamma\rightarrow\gamma\gamma$ process, where it also leads to an increase in the integrated cross sections due to higher statistics in the central region (Figure 3). For the other processes $\gamma\gamma\rightarrow\gamma Z$ and $\gamma\gamma\rightarrow ZZ$, the lower energy region is kinematically forbidden, and the total cross section for CAIN-based spectra is reduced compared to the LCA.
Figure 5: Final state transverse momentum kinematic distributions for different electron beam energies $\sqrt{s_{ee}}=500$ GeV (upper row) and 1000 GeV (lower row) and various polarizations (see text).
Another interesting observation on the Figures 4 and 5 is that the actual effect of the Compton non-linearity degree becomes more prominent at higher energies and is more noticeable on $p_{T}$ distributions. Such an effect is likely attributed to the fact that this observable is not Lorentz-invariant.
$n_{ph}$ 0 (LCA) 1 2
$\gamma\gamma\rightarrow\gamma\gamma$, CAIN 13.8457(7) 14.8800(7) 14.3350(7)
500 GeV SANCphot 8.0457(4) - -
$\gamma\gamma\rightarrow\gamma\gamma$, CAIN 10.3750(5) 8.9332(4) 9.3450(5)
1000 GeV SANCphot 6.8598(3) - -
$\gamma\gamma\rightarrow\gamma Z$, CAIN 8.7229(4) 8.8690(4) 8.7565(4)
500 GeV SANCphot 19.3739(1) - -
$\gamma\gamma\rightarrow\gamma Z$, CAIN 17.3690(9) 23.014(1) 21.136(1)
1000 GeV SANCphot 38.279(2) - -
$\gamma\gamma\rightarrow ZZ$, CAIN 9.2385(5) 8.3770(4) 8.5766(4)
500 GeV SANCphot 21.809(1) - -
$\gamma\gamma\rightarrow ZZ$, CAIN 25.644(1) 32.491(2) 30.055(1)
1000 GeV SANCphot 58.945(3) - -
Table 2: Integrated cross sections $\sigma(\gamma\gamma)$ [fb] with different orders $n_{ph}$ for linear polarization.
$n_{ph}$ 0 (LCA) 1 2
$\gamma\gamma\rightarrow\gamma\gamma$, CAIN 16.2857(8) 17.0891(9) 16.4330(8)
500 GeV SANCphot 9.4432(5) - -
$\gamma\gamma\rightarrow\gamma\gamma$, CAIN 13.2011(7) 10.6979(5) 11.6521(6)
1000 GeV SANCphot 6.9403(3) - -
$\gamma\gamma\rightarrow\gamma Z$, CAIN 8.2950(4) 9.0302(4) 8.5068(4)
500 GeV SANCphot 21.528(1) - -
$\gamma\gamma\rightarrow\gamma Z$, CAIN 13.6011(7) 18.2681(9) 16.1505(8)
1000 GeV SANCphot 35.158(2) - -
$\gamma\gamma\rightarrow ZZ$, CAIN 12.0139(6) 12.4726(6) 11.8503(6)
500 GeV SANCphot 32.335(2) - -
$\gamma\gamma\rightarrow ZZ$, CAIN 22.620(1) 30.058(1) 26.679(1)
1000 GeV SANCphot 59.679(3) - -
Table 3: Integrated cross sections $\sigma(\gamma\gamma)$ [fb] with different orders $n_{ph}$ for circular polarization.
From the total cross section values, we see that for the $\gamma\gamma\rightarrow\gamma\gamma$ process, adding a realistic beam geometry to the photon spectrum simulation increases the integrated cross section in the given calculation limits by a factor of 2 depending on whether the polarization is linear or circular. However, this effect is washed away for higher collision energies and other processes, where, due to the kinematic threshold, the effect becomes reversed, and for 500 GeV energy, the total cross section for the $\gamma\gamma\rightarrow\gamma Z$ process becomes 3 times less after the inclusion of non-linear effects and beam geometry.
Related to the BSM studies, it is also interesting to look at the separate cross section components, i.e. $\sigma_{22}$ and $\sigma_{3}$ which are sensitive to particular contributions of SUSY particles as stated in [13]. Figure 6 shows that adding a detailed beam geometry description leads to a dampening of these components, similar to the total cross section, without particular changes in the distribution shapes. The next orders of approximation exhibit shifting of the spectra to the lower energies.
Figure 6: Contributions of $\sigma_{22}$ and $\sigma_{3}$ to the kinematic distributions for $\gamma\gamma\rightarrow\gamma Z$ process at $\sqrt{s_{ee}}=500$ GeV. The line colors’ meaning is the same as in Figures 4 and 5
Being able to include a more realistic description of the photon spectra is definitely beneficial for precise studies of $\gamma\gamma$-induced processes. It would also be interesting to look at the higher orders of Compton non-linearity for linear polarizations, as this would allow us to test the different polarization variants depending on the $\phi$ direction. Another fruitful topic of future research would be adding the decay of the final Z particles and thus gaining access to a broad range of studies related to polar angle asymmetries. One more improvement to this study would be the inclusion of the depolarisation effects into account, which should further improve our precision.
As the input photon energy distribution is agnostic of the photon generation process, this new approach allows the simulation of the SANCphot processes with initial photons produced not only via Compton back-scattering, but also via the beamstrahlung [19, 26] given the spectra, generated, for example, with GuineaPig program [27].
6 Summary
Thanks to the inclusion of a piecewise description of general photon spectra along with the polarization information, the precision of photon-photon collision simulation in SANCphot is substantially improved. With realistic photon beam simulation by the CAIN program, we see that the total cross section of the $\gamma\gamma\rightarrow\gamma\gamma$ process is increased by 50 - 100% depending on the beam energy, and for the processes with final Z boson(s) is reduced by a factor of 2. In addition, the kinematic distributions show a noticeable shift to lower energies while keeping the polarization distinctive features intact.
7 Acknowledgements
We thank V. Telnov and I. Ginzburg for their advice and fruitful discussion.
Funding
This research has been funded by the Committee of Science of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP19680084).

### v2

@@
 While the phenomena of photon-photon collisions were previously observed as a side effect in electron-positron and heavy ion collisions, a dedicated machine would open up new avenues for precision electroweak physics.
-An overview presented in [1] covers a wide range of topics dedicated to the development of a photon-photon collider, including physics program, photon beam generation, and effects in the interaction region.
+An overview presented in [ECFADESYPhotonColliderWorkingGroup:2001ikq] covers a wide range of topics dedicated to the development of a photon-photon collider, including physics program, photon beam generation, and effects in the interaction region.
 Theoretical calculations and simulation codes accompany these developments, gradually improving our understanding and correcting expectations from these intriguing projects.
@@
 When studied in angular distributions, polarization effects of the Z-boson in processes $\gamma\gamma\rightarrow\gamma Z$ and $\gamma\gamma\rightarrow ZZ$ provide sensitivity to potential new physics beyond the Standard Model.
-Anomalous quartic gauge couplings and extra dimensions are another kind of phenomena, where accurate simulation of polarized photon-photon collisions can improve background estimation and lead to better constraints [2] in these domains.
-Previously, several remarkable works have been devoted to the study of $\gamma\gamma$ processes, in particular, light-by-light scattering (LbL), taking into account radiative corrections [3, 4, 5, 6, 7].
-The process of $\gamma\gamma\to\gamma Z$ was revized in [8].
-A calculation of ZZ production cross-section components ($\sigma_{0},\sigma_{22},$ etc.) was presented and available as a public code [9].
-Processes $\gamma\gamma\rightarrow\gamma\gamma$ and $\gamma\gamma\rightarrow\gamma Z$ with polarized incoming beams at CLIC are investigated in terms of anomalous couplings [10, 11].
+Anomalous quartic gauge couplings and extra dimensions are another kind of phenomena, where accurate simulation of polarized photon-photon collisions can improve background estimation and lead to better constraints [Brodsky:1994nf] in these domains.
+Previously, several remarkable works have been devoted to the study of $\gamma\gamma$ processes, in particular, light-by-light scattering (LbL), taking into account radiative corrections [Karplus:1950zz, Jikia:1993tc, Bohm:1994sf, Gounaris:1998qk, Bern:2001dg].
+The process of $\gamma\gamma\to\gamma Z$ was reviewed in [Gounaris:1999ux].
+A calculation of ZZ production cross-section components ($\sigma_{0},\sigma_{22},$ etc.) was presented and available as a public code [Diakonidis:2006cv].
+Processes $\gamma\gamma\rightarrow\gamma\gamma$ and $\gamma\gamma\rightarrow\gamma Z$ with polarized incoming beams at CLIC are investigated in terms of anomalous couplings [Inan:2020gfu, Inan:2021pbx].
 The paper is composed as follows.
 A short preliminary word on the current status of the photon-photon collider is given in the Introduction 1.
-We revise how photon-photon collisions are treated in the SANCphot code in Section 2.
+We review how photon-photon collisions are treated in the SANCphot code in Section 2.
 Then, in Section 3, we overview challenges occurring when trying to give a precise description of photon spectra generated at the photon-photon colliders by the Compton scattering.
@@
 2 Photon-photon collisions in SANC
-The cross section of photon-photon collisions implemented in the SANCphot tool [12] follows a recipe described in [13] using FORTRAN modules introduced in [14, 15, 16].
+The cross section of photon-photon collisions implemented in the SANCphot tool [Bondarenko:2022ddm] follows a recipe described in [Gounaris:1999gh] using FORTRAN modules introduced in [Bardin:2009gq, Bardin:2012mc, Bardin:2017qqw].
 The total $\gamma\gamma$ process cross section is represented as a convolution of its components in different helicity combinations $d\bar{\sigma}_{ij}/d\cos\vartheta^{*}$ and averaged Stokes parameters defining the density matrix of the back-scattered photon $\langle\xi_{i}\xi_{j}^{\prime}\rangle$:
-$\begin{split}{\frac{d\sigma}{d\tau d\cos\vartheta^{*}}}=&{\frac{d\bar{L}_{\gamma\gamma}}{d\tau}}\Bigg{\{}{\frac{d\bar{\sigma}_{0}}{d\cos\vartheta^{*}}}+\langle\xi_{2}\xi_{2}^{\prime}\rangle{\frac{d\bar{\sigma}_{22}}{d\cos\vartheta^{*}}}\\ &+[\langle\xi_{3}\rangle\cos 2\phi+\langle\xi_{3}^{\prime}\rangle\cos 2\phi^{\prime}]{\frac{d\bar{\sigma}_{3}}{d\cos\vartheta^{*}}}\\ &+\langle\xi_{3}\xi_{3}^{\prime}\rangle[{\frac{d\bar{\sigma}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi+\phi^{\prime})\\ &+{\frac{d\bar{\sigma}^{\prime}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi-\phi^{\prime})]\\ &+[\langle\xi_{2}\xi_{3}^{\prime}\rangle\sin 2\phi^{\prime}\\ &-\langle\xi_{3}\xi_{2}^{\prime}\rangle\sin 2\phi]{\frac{d\bar{\sigma}_{23}}{d\cos\vartheta^{*}}}\Bigg{\}}\ ,\end{split}$ (1)
-In this definition, the $\vartheta^{*}$ is the scattering angle of $\gamma\gamma$ collision products measured in the rest frame of the incoming particles, and the $\tau\equiv s_{\gamma\gamma}/s_{ee}$ — a fraction of the electron beam’s energy transferred to the back-scattered photons.
-Further, $d\bar{L}_{\gamma\gamma}/d\tau$ denotes the photon-photon luminosity per unit $e^{+}e^{-}$ flux, considering the linear collider operating in $\gamma\gamma$ mode, and the Stocks parameters $\xi_{2},\xi_{3}$ are defined via the photon and electron polarization functions as described in [17][18].
-The $d\bar{L}_{\gamma\gamma}/d\tau$ parameter, consequently, depends on the energy distribution of the back-scattered photons and can be evaluated analytically using the linear Compton approximation (LCA).
+$\begin{split}{\frac{d\sigma}{dx_{1}dx_{2}d\cos\vartheta^{*}}}=&{\frac{d\bar{L}_{\gamma\gamma}}{dx_{1}dx_{2}}}\Bigg\{{\frac{d\bar{\sigma}_{0}}{d\cos\vartheta^{*}}}+\langle\xi_{2}\xi_{2}^{\prime}\rangle{\frac{d\bar{\sigma}_{22}}{d\cos\vartheta^{*}}}\\ &+[\langle\xi_{3}\rangle\cos 2\phi+\langle\xi_{3}^{\prime}\rangle\cos 2\phi^{\prime}]{\frac{d\bar{\sigma}_{3}}{d\cos\vartheta^{*}}}\\ &+\langle\xi_{3}\xi_{3}^{\prime}\rangle[{\frac{d\bar{\sigma}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi+\phi^{\prime})\\ &+{\frac{d\bar{\sigma}^{\prime}_{33}}{d\cos\vartheta^{*}}}\cos 2(\phi-\phi^{\prime})]\\ &+[\langle\xi_{2}\xi_{3}^{\prime}\rangle\sin 2\phi^{\prime}\\ &-\langle\xi_{3}\xi_{2}^{\prime}\rangle\sin 2\phi]{\frac{d\bar{\sigma}_{23}}{d\cos\vartheta^{*}}}\Bigg\}\ ,\end{split}$ (1)
+In this definition, the $\vartheta^{*}$ is the scattering angle of $\gamma\gamma$ collision products measured in the rest frame of the incoming particles, and $x_{1},x_{2}$ — a fractions of the electron’s energy transferred to the back-scattered photon.
+Further, $\frac{d\bar{L}_{\gamma\gamma}}{dx_{1}dx_{2}}$ denotes the photon-photon luminosity per unit $e^{-}e^{-}$ flux, considering the linear collider operating in $\gamma\gamma$ mode, and the Stocks parameters $\xi_{2},\xi_{3}$ are defined via the photon and electron polarization functions as described in [Ginzburg:1981vm][Ginzburg:1982yr].
+The $\frac{d\bar{L}_{\gamma\gamma}}{dx_{1}dx_{2}}$ factor, consequently, depends on the energy distribution of the back-scattered photons and can be evaluated analytically using the linear Compton approximation (LCA).
 3 Photon spectrum in photon colliders
-The linear Compton approximation has several drawbacks when utilized in the simulation of the $\gamma\gamma$ collisions in the photon-photon collider [19].
+The linear Compton approximation has several drawbacks when utilized in the simulation of the $\gamma\gamma$ collisions in the photon-photon collider [Telnov:1989sd].
 Most of the inaccuracy is attributed to the process of electron beam conversion to a high-energy photonic beam.
@@
 Several works have been attributed to improving the spectrum of the photons generated via electron energy transfer.
-To our knowledge, the said issues are best addressed in the photon collider simulation code CAIN [20].
+To our knowledge, the said issues are best addressed in the photon collider simulation code CAIN [Chen:1994jt].
 The substantial improvements are achieved by introducing full treatment of nonlinear and polarized Compton scattering, a detailed description of the space-time structure of the electron and laser beams, complete kinematics, including angular distributions, and interplay with beamstrahlung and other background processes.
-Other efforts to improve photon spectrum approximation are presented in works [21, 22] which contribute to a detailed description of nonlinear effects and a complete description of the photon polarization in Compton back-scattering.
-Public codes suitable for photon spectrum simulation include CompAZ [23] and CIRCE1 [24], both based on the creation of explicitly parameterized distributions.
+Other efforts to improve photon spectrum approximation are presented in works [Galynskii:2000fk, Ivanov:2004fi] which contribute to a detailed description of nonlinear effects and a complete description of the photon polarization in Compton back-scattering.
+Public codes suitable for photon spectrum simulation include CompAZ [Zarnecki:2002qr] and CIRCE1 [Ohl:1996fi], both based on the creation of explicitly parameterized distributions.
 These tools are focused on approximating the high-energy regions of the photon spectra.
-CIRCE2 [25] features correlated double-beam spectra based on tabulated energy distributions built from the event records generated in advance.
+CIRCE2 [Ohl:2002ff] features correlated double-beam spectra based on tabulated energy distributions built from the event records generated in advance.
 4 Coupling CAIN and SANCphot
-In order to supply the realistic photon spectra obtained with the CAIN simulation, including the polarization information, to the Monte Carlo integration routines implemented in SANCphot, we use an approach similar to CIRCE2 [25].
+In order to supply the realistic photon spectra obtained with the CAIN simulation, including the polarization information, to the Monte Carlo integration routines implemented in SANCphot, we use an approach similar to CIRCE2 [Ohl:2002ff].
 The event records, generated with a realistic laser and oncoming electron beam parameters, along with a non-linear Compton process expressions, are aggregated into piecewise continuous functions, representing approximate distributions of spectra and linear and circular polarizations.
-$f_{L}(x)=\sum_{i=0}^{n}\bar{L}_{i}\chi_{i}(x),\;\chi_{i}(x)=\begin{cases}1\;\mathrm{if}\;\left\llbracket x/n\right\rrbracket=1\\ 0\;\mathrm{if}\;\left\llbracket x/n\right\rrbracket\neq 1\\ \end{cases}$ (2)
+[…]
 where $f_{L}$ is a piecewise approximation of the luminosity, $\chi_{i}(x)$ is an indicator function, $L_{i}$ is the aggregated $L_{\gamma\gamma}$ value in bin $i$.
@@
 The cubic interpolation of these functions serves as an input spectrum to SANCphot’s integration procedures in the same manner as analytic expressions of linear Compton.
-As an example of a realistic photon spectrum simulation, the parameters of the CLIC accelerator was chosen to configure the CAIN code: number of electrons per bunch $N=4\times 10^{9}$, the RMS (root-mean-square) bunch length in the longitudinal direction $\sigma_{z}=0.05$ mm, normalized emittance in the y direction $\gamma\varepsilon_{y}=1\times 10^{-5}\mathrm{m}\cdot\mathrm{rad}$ and in x direction $1.9\times 10^{-6}$ and $1.5\times 10^{-6}\mathrm{m}\cdot\mathrm{rad}$ for 500 and 1000 GeV respectively, beta functions $\beta_{x,y}=2/0.1$ mm and the transverse beam sizes $\sigma_{x,y}=88/4.5$ nm for 500 GeV and $55/3.2$ nm for 1000 GeV.
+[…]
 The polarizations of the back-scattered photons in CAIN were controlled via Stokes parameters $(\xi_{1},\xi_{2},\xi_{3})$.
@@
 The difference between $n_{ph}=0$ and $n_{ph}=1$ setting is that the former uses linear Compton formulas, which are inapplicable in strong laser fields.
-In the CAIN simulation, the strength of a laser electromagnetic field is characterized by a parameter $\xi=\frac{\lambda_{L}}{m}\sqrt{\mu_{0}cP}$, where $\lambda_{L}$ is the laser wavelength, m is the electron mass, $c$ is the speed of light, $\mu_{0}=4\pi\times 10^{-7}$, and P is the laser power density.
+The strength of a laser electromagnetic field is characterized by a parameter $\xi^{2}=2n_{\gamma}r_{e}^{2}\lambda/\alpha$, where $n_{\gamma}$ is the density of laser photons, $\lambda$ is the laser wavelength, $\alpha=e^{2}/\hbar c=1/137$, $r_{e}$ - electron radius [ECFADESYPhotonColliderWorkingGroup:2001ikq, Ivanov:2004fi].
 At $\xi<O(1)$, the non-linear Compton expression requires expansion in terms of $n_{ph}$ to provide a better description.
-Figure 1: A comparison of photon energy spectra $dL/dx$ and cross section polarizing coefficients $d\langle\xi 2\rangle/dx$ and $d\langle\xi 3\rangle/dx$ depending on the fraction of transferred energy $x$.
-Solid lines and dashed lines correspond to linear and circular laser polarizations.
+For the laser configuration provided in the above paragraph $\xi^{2}\approx 0.3$ and the number of photons per electron is $k=4.1$ for 500 GeV and $5.6$ for 1000 GeV, which justifies modeling the process as multiple Compton scattering.
+Figure 1: A comparison of normalized photon energy spectra $\frac{1}{N_{\gamma}}\frac{dN_{\gamma}}{dx}$ and cross section polarizing coefficients $d\langle\xi 2\rangle/dx$ and $d\langle\xi 3\rangle/dx$ depending on the fraction of transferred energy $x$.
+The left and right columns correspond to linear and circular laser polarizations respectively.
 In the present study, we use two distinct polarization configurations:
@@
 Figure 1 shows photon energy spectra and polarization at different levels of approximation: LCA calculated with SANCphot (black lines), linear and non-linear Compton approximations calculated by CAIN with $n_{ph}=$ 0 (red), 1 (blue), and 2 (green) with realistic beam geometry.
-The CAIN spectra show a significant shift towards lower energy regions due to multiple electron-photon interactions and spatial beam effects.
+The red curve shows effects related to the multiple photon scattered per electron, the blue curves depict engagement of non-linear calculations, giving significant shift towards lower energy regions, but the green curves corresponding to higher harmonics show slight shift towards higher energy regions which complies with Fig.2 of [Ivanov:2004fi].
 5 Numerical results
-The numerical validation was performed in the $\alpha(0)$ EW scheme using the setup given below. For ease of validation, we keep the electroweak parameter values the same as in [12]:
+The numerical validation was performed in the $\alpha(0)$ EW scheme using the setup given below. For ease of validation, we keep the electroweak parameter values the same as in [Bondarenko:2022ddm]:
 $\alpha=1/137.035990996$, $G_{F}=1.13024\times 10^{-5}\,\mathrm{GeV}^{-2}$, $M_{W}=80.45149$ GeV, $M_{Z}=91.18670$ GeV, $M_{H}=125$ GeV, $m_{e}=0.51099907$ MeV, $m_{\mu}=0.105658389$ GeV, $m_{\tau}=1.77705$ GeV, $m_{u}=0.062$ GeV, $m_{d}=0.083$ GeV, $m_{c}=1.5$ GeV, $m_{s}=0.215$ GeV, $m_{t}=173.8$ GeV, $m_{b}=4.7$ GeV.
-The simulation of gamma-gamma collisions was performed as if the photons were produced in the $ee$ collider framework.
 Initial photon spectra were approximated using several methods: (i) analytic linear Compton approximation implemented in SANCphot, (ii) photon energy distribution representing linear Compton spectrum, (iii) realistic spectra obtained from CAIN photon collider simulation package.
@@
 Lower plots give the ratios of distributions built using piecewise approximation of SANCphot (red) and CAIN (blue) spectra to the analytic values.
-In addition, Table 1 shows numeric comparison of three different sources of photon spectra, when using the analytic version from SANCphot, piecewise spectrum description of the analytic expression, and LCA obtained with CAIN minimizing beam geometry effects (i.e., using a very large beam, very low laser intensity, and very low number of particles per bunch).
+In addition, Table 1 shows numeric comparison of three different sources of photon spectra, when using the analytic version from SANCphot, piecewise spectrum description of the analytic expression, and LCA obtained with CAIN minimizing beam geometry effects.
+For the latter the laser parameters were set to $z_{R}=1$ $m$, $\sigma_{t}=100$ ps, which corresponds to $\xi^{2}\approx 3\times 10^{-4}$, ensuring the linear mode for Compton back-scattering.
 We see that the piecewise description of pure LCA spectrum is very close to the analytic one.
@@
 Figure 3: Final state $\cos\theta$ kinematic distributions for electron beam energy $\sqrt{s_{ee}}=500$ GeV and various polarizations (see text).
+The color legend for $\gamma\gamma\rightarrow\gamma\gamma$ plot applies to the other two.
 Figure 4: Final state invariant mass kinematic distributions for different electron beam energies $\sqrt{s_{ee}}=500$ GeV (upper row) and 1000 GeV (lower row) and various polarizations (see text).
+The color legend for 500 GeV $\gamma\gamma\rightarrow\gamma\gamma$ plot applies to the others.
 In the invariant mass distributions, we observe a well-known energy-dampening effect of taking into account realistic beam geometry.
@@
 Figure 5: Final state transverse momentum kinematic distributions for different electron beam energies $\sqrt{s_{ee}}=500$ GeV (upper row) and 1000 GeV (lower row) and various polarizations (see text).
+The color legend for 500 GeV $\gamma\gamma\rightarrow\gamma\gamma$ plot applies to the others.
 Another interesting observation on the Figures 4 and 5 is that the actual effect of the Compton non-linearity degree becomes more prominent at higher energies and is more noticeable on $p_{T}$ distributions.
@@
 Related to the BSM studies, it is also interesting to look at the separate cross section components, i.e.
-$\sigma_{22}$ and $\sigma_{3}$ which are sensitive to particular contributions of SUSY particles as stated in [13].
+$\sigma_{22}$ and $\sigma_{3}$ which are sensitive to particular contributions of SUSY particles as stated in [Gounaris:1999gh].
 Figure 6 shows that adding a detailed beam geometry description leads to a dampening of these components, similar to the total cross section, without particular changes in the distribution shapes.
@@
 Figure 6: Contributions of $\sigma_{22}$ and $\sigma_{3}$ to the kinematic distributions for $\gamma\gamma\rightarrow\gamma Z$ process at $\sqrt{s_{ee}}=500$ GeV.
-The line colors’ meaning is the same as in Figures 4 and 5
+The line colors’ meaning is the same as in Figures 4 and 5.
 Being able to include a more realistic description of the photon spectra is definitely beneficial for precise studies of $\gamma\gamma$-induced processes.
@@
 One more improvement to this study would be the inclusion of the depolarisation effects into account, which should further improve our precision.
-As the input photon energy distribution is agnostic of the photon generation process, this new approach allows the simulation of the SANCphot processes with initial photons produced not only via Compton back-scattering, but also via the beamstrahlung [19, 26] given the spectra, generated, for example, with GuineaPig program [27].
+As the input photon energy distribution is agnostic of the photon generation process, this new approach allows the simulation of the SANCphot processes with initial photons produced not only via Compton back-scattering, but also via the beamstrahlung [Telnov:1989sd, Yokoya:1991qz] given the spectra, generated, for example, with GuineaPig program [Schulte:1998au].
 6 Summary